OpenStudy (anonymous):

find formulae for fogoh f(x)=1/1+x , g(x)=3root x , h(x)=1/x power 3

hartnn (hartnn):

first find goh (x),can u ?

OpenStudy (anonymous):

no

OpenStudy (mertsj):

goh means replace each x in the function g with 1/x

OpenStudy (mertsj):

So do that and post your result.

hartnn (hartnn):

goh (x) can be found by replacing 'x' in g(x) by h(x)=1/x^3

OpenStudy (anonymous):

f{g(1/x^3)}

OpenStudy (mertsj):

Replace x in the function f with that result and you will be done except for simplification

hartnn (hartnn):

see this : when i replace 'x' in g(x) by 1/x^3, i get goh(x) = \(3 \sqrt{1/x^3}=3 \sqrt{x^{-3}}\) got this ?

OpenStudy (anonymous):

can i show my whole solution then you tell me

hartnn (hartnn):

sure, if incorrect, we'll spot the error.

OpenStudy (anonymous):

ok wait for a while

OpenStudy (anonymous):

f [ g {h(x) } ] f [ g (1/x^3) ] f [ 3root 1/x^3 ] f [ 3 root x power -3 ] tell me is this ans ?

hartnn (hartnn):

thats just \(f(3\sqrt{x^{-3}})\) you need to evaluate it by replacing x by \(3\sqrt{x^{-3}}in f(x )\)

OpenStudy (mertsj):